OpenStudy (anonymous):

Help would be really appreciated! Find all the missing parts to the triangle below...

OpenStudy (ria23):

Do yhu know the SAS formula?

OpenStudy (anonymous):

Yes, but the problem is related to the Law of Cosines.

OpenStudy (anonymous):

@ranga

OpenStudy (ria23):

Ok. c:

OpenStudy (ria23):

Yhur formula is Area= 1/2 ab sin (C)

OpenStudy (anonymous):

Are you sure..?

OpenStudy (dumbcow):

she doen't need area?

OpenStudy (ria23):

Yes, Do yhu know how to work that out?

OpenStudy (ria23):

No, Yhu have to find the missing parts of the triangle, the missing measurements.

OpenStudy (anonymous):

I only need the missing leg and angles.

OpenStudy (ria23):

Using that formula, yhu'll get yhur missing leg.

OpenStudy (ria23):

Oh wait... Hang on.

OpenStudy (dumbcow):

law of cosines \[c^2 = a^2 + b^2 - 2ab \cos C\] law of sines \[\frac{\sin A}{a} = \frac{\sin C}{c}\]

OpenStudy (ria23):

Beat me to it. :D

OpenStudy (dumbcow):

a = 15 b = 28 C = 87

OpenStudy (anonymous):

c^2=15^2+28^2-2(15)(28)cos87

OpenStudy (anonymous):

Right?

OpenStudy (ria23):

Yes.

OpenStudy (anonymous):

I got 965.03

OpenStudy (anonymous):

√965.03=31.06 , correct?

OpenStudy (ria23):

I got 225+784-840= 169 times cos87= 8.84

OpenStudy (anonymous):

@mathmale

OpenStudy (ria23):

@tgawade assistance is required.

OpenStudy (dumbcow):

@Cubi-Cal , you are correct c = 31.065

OpenStudy (mathmale):

D. C. provided the correct formula. D. C.: Please discuss with Cubi-Cal how you arrived at this final result, so that all the steps of the process are clear. Thanks.

OpenStudy (anonymous):

Okay great! So that would take care of which part of the triangle?

OpenStudy (dumbcow):

the 3rd leg...part "f"

OpenStudy (anonymous):

Ah okay.

OpenStudy (anonymous):

So how do I use f to solve the angles, now?

OpenStudy (anonymous):

since (GH)^2 = 980.73 so GH = 31.316

OpenStudy (anonymous):

Use laws of sine

OpenStudy (dumbcow):

Law of sines \[\frac{\sin H}{15} = \frac{\sin 87}{31.065}\]

OpenStudy (anonymous):

SinH31.065=14.97

OpenStudy (anonymous):

Correct?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

Okay, then would I do arcsin14.97? Which gave me .99

OpenStudy (mathmale):

I'm assuming that this is the problem you wanted to discuss. Is it? If so, let's summarize what we already know: 2 sides are 15 and 28, and we've found f, the third side. What is it? We have one angle, which is 87 degrees. Correct?

OpenStudy (anonymous):

Yes this is the problem! f=31.065

OpenStudy (mathmale):

Which of the unknown angles would you like to find first? The known angle is 87 degrees.

OpenStudy (anonymous):

We could do G first.

OpenStudy (mathmale):

Have you used the Law of Sines before?

OpenStudy (anonymous):

Yes!

OpenStudy (mathmale):

Great. You want angle G. You have side g = ??

OpenStudy (anonymous):

g would be 15, right?

OpenStudy (mathmale):

It's side h that is 15. Note that side h is opposite angle H. You want angle G. the side opposite angle G is g = ??

OpenStudy (mathmale):

Think: g is opposite angle G, h is opposite angle H.

OpenStudy (anonymous):

28!

OpenStudy (anonymous):

To find angle G use laws of sine ... sinG/28 = sin87/f

OpenStudy (mathmale):

Right. You want angle G. You already have side g, which is 28, as well as h, which is 15, and f, which is 31.065. Write out this equation: sin G sin 87 ---- = ---------- g 31.065

OpenStudy (mathmale):

This is the Law of Sinces applied to finding angle G. Actually, we find sin G first, then G itself. Want G in degrees or rad?

OpenStudy (anonymous):

SinG31.065=.99g

OpenStudy (anonymous):

Probably rad.

OpenStudy (mathmale):

(Looking at your equation.) But since angle F is given in degrees (87 degrees), you should look for angle G in degrees also.

OpenStudy (anonymous):

Oh right.

OpenStudy (mathmale):

Strong suggestion: Use parentheses to make your equation easier to read and understand: 31.065sin G = (g)sin 87

OpenStudy (mathmale):

But g is known, isn't it? If so, replace g with its value in your equation, above.

OpenStudy (anonymous):

SinG 31.065=27.96

OpenStudy (mathmale):

That looks good; I'd been looking for 31.065 sin G = 28 sin 87.

OpenStudy (mathmale):

How would you now isolate "sin G"?

OpenStudy (anonymous):

By using arcsin?

OpenStudy (mathmale):

not yet; have to have sin G alone on the left side first. After that, then yes, you do need to use the arcsin function.

OpenStudy (anonymous):

Um.. division then?

OpenStudy (anonymous):

If not subtraction?

OpenStudy (mathmale):

I'd prefer you write this equation as 31.065 sin G = 28 sin 87. Divide both sides by 31.065 now. No subtraction is involved here.

OpenStudy (anonymous):

Okay I got .9001

OpenStudy (mathmale):

So did I! Now please find the inverse sine of that.

OpenStudy (anonymous):

Okay, great! That would be 64.17

OpenStudy (mathmale):

When I round that off, I get 64.17. Actually, with quantities given in the original problem as 15, 87 and 28, we should be rounding everythign off to the nearest integer. But for now, say G = 64.2 degrees. Agree with that? You wanted G.

OpenStudy (anonymous):

Got it! I agree!

OpenStudy (anonymous):

So lastly is H.

OpenStudy (mathmale):

Now you have a triangle with two angles known: they are 64 and 87. Hint: What is the sum of the interior angles of a triangle? How would u use this fact to find H?

OpenStudy (anonymous):

OH! 64+87+29=180

OpenStudy (mathmale):

Cool.

OpenStudy (mathmale):

any questions about using either the Law of Sines or the Law of Cosines?

OpenStudy (anonymous):

Yes, would you mind checking my work in a previous question?

OpenStudy (mathmale):

Could you possibly give me the URL of that question?

OpenStudy (anonymous):

Yes.. http://openstudy.com/study#/updates/4f887586e4b02251ecc96433